Suppose C&Y restaurant borrow a 5-year loan of $85,000 at an annual interest rate of 5%. The loan agreement states that the repayment of principal and the loan interest has to be paid by the end of each year. The instalment of each repayment is fixed amount throughout the loan period. You are instructed to construct an amortization schedule for loan repayment including beginning balance, annual payment, interest and ending balance.
Answer #1

The first step is to calculate the annual payment on the loan. The following equation is used to calculate the value of the annual payment.

PMT Present value of the loan 1- (1 + interest rate)-nt interest rate

1590396663911_blob.png

The annual payment on the loan equals $ 19632.86 \approx $ 19632.9

The amortization schedule for the loan payment is done using excel.

Year Beginning balance ($) Annual payment ($) Interest ($) Principal paid ($) Ending balance ($)
1 85000 19632.9 0.05 × 85000 = 4250 19632.9 - 4250 = 15382.9 85000 - 15382.9 = 69617.1
2 69617.1 19632.9 0.05 × 69617.1 = 3480.9 19632.9 - 3480.9 = 16152 69617.1 - 16152 = 53465.1
3 53465.1 19632.9 0.05 × 53465.1 = 2673.3 19632.9 - 2673.3 = 16959.6 53465.1 - 16959.6 = 36505.5
4 36505.5 19632.9 0.05 × 36505.5 = 1825.3 19632.9 - 1825.3 = 17807.6 36505.5 - 17807.6 = 18697.9
5 18697.9 19632.9 0.05 × 18697.9 = 934.9 19632.9 - 934.9 = 18698 18697.9 - 18698   ̴ 0